### v1

# Phenomenological implications of a class of non-invertible selection rules

Abstract
Through well-motivated models in particle physics, we demonstrate the power of a general class of selection rules from non-invertible fusion algebra that are only exact at low orders in perturbation theory, recently introduced in Kaidi et al. (2024). Surprisingly, these non-invertible selection rules can even be applied to the minimal extension of the Standard Model, which is to add a gauge-singlet real scalar. In this model, we show that Fibonacci fusion rules can lead to experimentally testable features for the scattering processes of the real scalar. We anticipate that this class of non-invertible selection rules can be applied to a wide range of models beyond the Standard Model. To further strengthen our methodology, we discuss a dark matter model based on the Ising fusion rules, where the dark matter is labeled by the non-invertible element in the algebra, hence its stability is preserved at all loop orders.
I Introduction
The notion of symmetries has been generalized significantly in the past decade since the seminal work Gaiotto et al. (2015). Perhaps the most drastic generalization is the so-called non-invertible symmetries; see e.g. Cordova et al. (2022); Shao (2023); Schafer-Nameki (2024); Costa et al. (2024); Brennan and Hong (2023) for reviews. Conventionally, these symmetries appear in two-dimensional rational conformal field theories Chang et al. (2019); Thorngren and Wang (2024a, b), and their generalizations in four-dimensional quantum field theories Choi et al. (2022a); Kaidi et al. (2022) and applications in particle physics Choi et al. (2022b); Cordova and Ohmori (2023); Putrov and Wang (2024); Choi et al. (2024); Cordova et al. (2024a, b, c); Delgado and Koren (2025); Kobayashi et al. (2024); Funakoshi et al. (2024); Kobayashi et al. (2025); Cao et al. (2024) has just been uncovered in recent years. For the interest of particle physics, celebrated examples of non-invertible symmetries include the Adler-Bell-Jackiw anomaly in quantum electrodynamics Choi et al. (2022b); Cordova and Ohmori (2023) and the higher-form generalizations in axion-Maxwell theory Choi et al. (2023); Yokokura (2022); Hidaka et al. (2024); Del Zotto et al. (2024).
More broadly, the existence of non-invertible symmetries suggests an intriguing possibility that the selection rules do not have to obey the group laws when they act on the fields. In this paper, we are interested in a general class of selection rules induced by non-invertible fusion algebra that are only exact in low orders in perturbation theory Kaidi et al. (2024). We call them non-invertible selection rules. We aim to show their usefulness in uncovering some peculiar features in particle physics models, which would otherwise be difficult to explain using the selection rules imposed by ordinary symmetries. Surprisingly, these non-invertible selection rules are even useful in the simplest extensions of the Standard Model (SM). Throughout the paper, we assume that the new particles are weakly coupled, therefore perturbation theory can apply.
Let us consider a finite set $A=\{\mathbbm{1},x,y,z,\cdots\}$ of basis elements with the fusion rules $x\otimes y=\sum_{z\in A}N^{z}_{x,y}z$, where $\mathbbm{1}$ is the identity element and $N^{z}_{x,y}$ are the structure constants. In general, the fusion rules are non-invertible. We will only consider the commutative fusion rules, i.e. $x\otimes y=y\otimes x$. One can construct a theory as follows Kaidi et al. (2024):
  • Each field in the theory $\phi_{i}$ is labeled by an element in the set, whose multiplications obey the fusion rules.
  • We assume the existence of the other element which labels the conjugated field $\bar{\phi}_{i}$, such that the identity element $\mathbbm{1}$ exists in the fusion product $\phi_{i}\otimes\bar{\phi}_{i}$.
  • A bare interaction term in the classical Lagrangian is allowed when the identity element is contained in the fusion product of the corresponding particles in the interaction.
As such, one constructs all the possible operators consistent with the fusion rules, and they correspond to the tree-level scattering amplitudes. Notice that the kinetic terms of $\phi_{i}$ are always allowed in the above construction. This needs to be contrasted with the background fields which do not propagate, e.g. one can promote coupling constants into the background fields in spurion analysis.
On the other hand, the operators that violate this class of non-invertible selection rules can be generated in radiative corrections as the loop order is increased. Hence these non-invertible selection rules are only exact at low orders in perturbation theory, at least tree-level exact. Naively, this contradicts the intuition of spurion analysis for the selection rules imposed by the ordinary symmetries, where the selection rules cannot be violated at the loop level in perturbation theory if they are exact at the tree level. However, due to the non-invertible nature, this kind of intuition is invalid for the selection rules discussed in this paper.
In the rest of the paper, we will demonstrate these features in two models in particle physics, arguably the simplest extensions of the SM.
II A real scalar and Fibonacci fusion rules
Let us start by adding a gauge-singlet real scalar to the SM. Despite its simplicity, the model is well-motivated by the rich phenomenology O’Connell et al. (2007); Chen et al. (2015); Buttazzo et al. (2015); He and Zhu (2017); Cao et al. (2018); Corbett et al. (2018); Li et al. (2019); Haisch et al. (2020); Adhikari et al. (2021); Dawson et al. (2021); Lewis et al. (2024). We denote the new scalar particle as $s$, it can have self-interactions and interactions with the Higgs doublet $H$. In particular, we consider the following Lagrangian at the classical level,
$\mathcal{L}=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}s)^{2}-\frac{m_{s}^{2}}{2}s^{2}-\frac{\lambda_{3}}{3!}s^{3}-\frac{\lambda_{4}}{4!}s^{4}-\frac{\lambda_{sh}}{2}s^{2}|H|^{2}\;.$ (1)
Such a Lagrangian is an unconventional one from the perspective of the ordinary $\mathbb{Z}_{2}$ symmetry, under which $s\to-s$ while all the SM particles are uncharged. If the $\mathbb{Z}_{2}$ symmetry is imposed, all the terms with odd powers of $s$ are forbidden, and the $\mathbb{Z}_{2}$ symmetry is exact at all loop orders in perturbation theory. However, the existence of $\lambda_{3}s^{3}$ implies that the $\mathbb{Z}_{2}$ symmetry is explicitly broken, hence there should also be terms such as $\lambda^{\prime}_{sh}s|H|^{2}$ in Eq. (1), if the theory is not fine-tuned. Since there is no obvious “symmetry” that can explain the coupling structure in Eq. (1), it is considered unnatural  [^1] Roughly speaking, the usual naturalness criterion states that a parameter in the Lagrangian is naturally small if there is an enhanced symmetry in the limit when the parameter vanishes ’t Hooft (1980). that $\lambda_{3}s^{3}$ is present in the classical Lagrangian while $\lambda^{\prime}_{sh}s|H|^{2}$ is not. The three-point vertex corresponding to the coupling $\lambda^{\prime}_{sh}$ is nevertheless generated at one-loop level when $\lambda_{3}$ and $\lambda_{sh}$ are present at tree level, i.e.
$(\lambda^{\prime}_{sh})^{\text{1-loop}}\sim\frac{\lambda^{\text{tree}}_{3}\lambda^{\text{tree}}_{sh}}{16\pi^{2}}\ .$ (2)
This equation can be interpreted as the relation between the scattering amplitudes of the processes $s\to H^{\dagger}H$, $s\to ss$, and $ss\to H^{\dagger}H$  [^2] Notice that some particles have to off-shell due to kinematics. However, our focus here is the dynamics, namely the interactions between particles. The same consideration applies to the models in the next section., where the first amplitude is obtained by gluing together the legs of $s$ in the second and the third amplitudes.
Based on the above discussion, a natural question arises:
Is there a “symmetry” (or a set of selection rules) that allows the non-vanishing bare couplings $\lambda_{3},\lambda_{sh},m_{s}^{2},\lambda_{4}$ but forbids $\lambda^{\prime}_{sh}$ in the classical Lagrangian?
If such a pattern is realized by a “symmetry”, we understand that the scattering amplitude of $s\to H^{\dagger}H$ can naturally be suppressed by a loop factor compared to that of $s\to ss$ and $ss\to H^{\dagger}H$.
We propose to use the class of non-invertible selection rules introduced in Kaidi et al. (2024) to address the above question. One peculiar feature is that these selection rules are tree-level exact but will be violated at loop orders in radiative corrections. Hence they are useful in explaining the hierarchies between couplings with the suppression given by the loop factors. Specifically, we use the Fibonacci fusion rules  [^3] Conventionally, the Fibonacci fusion rules are used to describe anyons Trebst et al. (2008) in condensed matter physics. Recently, the same algebra also appeared in flavor models explaining the Yukawa texture zeros based on $\mathbb{Z}_{2}$ gauging of a $\mathbb{Z}_{3}$ symmetry Kobayashi et al. (2024); Funakoshi et al. (2024); Kobayashi et al. (2025)., whose algebra is given by
$\mathbbm{1}\otimes\mathbbm{1}=\mathbbm{1},\quad\quad\mathbbm{1}\otimes\tau=\tau\otimes\mathbbm{1}=\tau,\quad\quad\tau\otimes\tau=\mathbbm{1}\oplus\tau\;.$ (3)
Notice that $\tau$ does not have the inverse, hence this algebra is non-invertible.  [^4] Intuitively, this algebra is very similar to an ordinary $\mathbb{Z}_{2}$ symmetry, that is why it can be relevant to this real scalar model. However, the crucial difference is only due to the non-invertible nature of $\tau$. As we will justify, this is crucial to forbid all the tree-level scattering amplitudes linear in the scalar $s$. In the model, we label the real scalar $s$ by the element $\tau$ and all the other SM particles by the identity $\mathbbm{1}$. It is easy to check that all the terms in Eq. (1) are consistent with Fibonacci fusion rules, namely the identity element is produced in the fusion product of the corresponding particles in each of the interactions. More importantly, all of the terms (up to dimension four) consistent with Eq. (3) are already included in the classical Lagrangian. Therefore, the Fibonacci fusion rules are tree-level exact, and they forbid the tree scattering process $s\to H^{\dagger}H$ in a fashion consistent with the usual naturalness criterion (rather than fine-tuning). In contrast, the scattering processes $s\to ss$ and $ss\to H^{\dagger}H$ are not forbidden at the tree level.
More generally, the selection rules in Eq. (3) imply that all the single-$s$ scattering processes must be loop induced, while the multi-$s$ scatterings can be present at tree-level; see in Figure. 1. Indeed, the operators of the type $s\cdot\mathcal{O}_{\text{SM}}$ are forbidden by Eq. (3), hence the corresponding scattering amplitudes must be induced at loop orders. On the other hand, operators of the types $s^{n}\cdot\mathcal{O}_{\text{SM}}$ are consistent with the selection rules in Eq. (3) for any $n>1$ and $n\in\mathbb{Z}$, hence they are allowed at tree level. Here $\mathcal{O}_{\text{SM}}$ denotes any operator consisting of only the SM particles. Since all the SM particles are labeled by the identity element, the Fibonacci fusion rules do not constrain the interactions in $\mathcal{O}_{\text{SM}}$.
Figure 1: A cartoon showing the scattering processes of the real scalar $s$ with the constraints imposed by the Fibonacci fusion rules, where all the single-$s$ scatterings must be loop induced (denoted by the gray blob), while the multi-$s$ scatterings can be present at either tree or loop levels (denoted by the white blob) in perturbation theory. Hence, if the real scalar is discovered by future experiments, one can potentially test the Fibonacci fusion rules in these scattering processes.
So far we have applied the Fibonacci fusion rules in Eq. (3) to the real scalar model, it follows that the scatterings of the scalar $s$ need to satisfy the relations such as in Eq. (2). A more general interpretation of this result is that all the single-$s$ scattering processes must be loop-induced. It would be ideal to understand Eq. (2) from the perspective of spurion analysis, which requires to promote the coupling constants into background fields labeled by either $\mathbbm{1}$ or $\tau$, such that the Fibonacci fusion rules are formally preserved. In this example, we see that the tree couplings $\lambda_{3},\lambda_{sh}$ can be labeled by either $\mathbbm{1}$ or $\tau$, but the loop-generated coupling $\lambda^{\prime}_{sh}$ must be labeled by $\tau$. Furthermore, to make Eq. (2) consistent with Eq. (3), at least one of the tree couplings $\lambda_{3},\lambda_{sh}$ should be labeled by $\tau$. We remark that this is a feature rather distinct from the spurion analysis for ordinary symmetries, where there is no obstruction to label all the tree couplings as identity if the symmetry is tree-level exact. A full understanding of the spurion analysis for this class of non-invertible selection rules will be pursued elsewhere.
It is somewhat unexpected that the non-invertible fusion rules in Eq. (3) can give a unified description of the Fibonacci sequence $\{0,1,1,2,3,5,8,13,21,\cdots\}$, where these numbers count how many times the identity element appears in each step of the fusion of $\tau$ with itself, and the scattering properties of a real scalar with all the SM particles. We view this as the simplest example showing the power of the non-invertible selection rules in particle physics.  [^5] It may seem overkill to apply the non-invertible selection rules to such a simple and calculable model. However, we find this example useful to streamline the overall logic. In a similar spirit, another example of this kind includes analyzing the ’t Hooft anomaly in the solvable quantum mechanics model Gaiotto et al. (2017); Cheng and Seiberg (2023).
III Dark matter and Ising fusion rules
Next, we apply the non-invertible selection rules to dark matter (DM) models, showing that our methodology can potentially have a wide range of applications.
Let us introduce a gauge-singlet real scalar $s$ and a gauge-singlet Majorana fermion $\chi$. They can have interactions described by the classical Lagrangian as follows,
$\mathcal{L}$ $=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}s)^{2}-\frac{m_{s}^{2}}{2}s^{2}-\frac{\lambda_{s}}{4!}s^{4}-\frac{1}{2}\lambda_{sh}s^{2}|H|^{2}$ (4)
[…]
In particular, we assume that the terms $s^{3}$ and $s|H|^{2}$ are not allowed in the classical Lagrangian. It follows that the scattering processes $s\to ss$ and $s\to H^{\dagger}H$ must be loop-induced in radiative corrections. For instance, $s\to ss$ can be induced at the one-loop level by a triangle loop of $\chi$ due to the tree amplitude $s\to\chi\chi$. In turn, $s\to H^{\dagger}H$ can be induced at the two-loop level by gluing together the legs of $s$ in the one-loop amplitude $s\to ss$ and the tree amplitude $ss\to H^{\dagger}H$.
Let us analyze the “symmetry” structure of the classical Lagrangian in Eq. (4). In the paradigm of ordinary symmetries, it is difficult to explain the following two features simultaneously.
  1. 1.
    It is not possible to assign a $\mathbb{Z}_{2}$-odd charge to the scalar $s$ due to the presence of the $s\chi^{2}$ term in Eq. (4), despite the fact that all the other couplings of $s$ respect a $\mathbb{Z}_{2}$ symmetry where $s\to-s$. Overall, since there is no $\mathbb{Z}_{2}$ symmetry protecting the interactions of $s$, the absence of the terms $s^{3}$ and $s|H|^{2}$ in the classical Lagrangian suggests fine-tuning of the corresponding coupling constants, naively.
  2. 2.
    However, we notice that Eq. (4) enjoys an invertible $\mathbb{Z}_{2}$ symmetry under which $\chi\to-\chi$ while all the other particles are charged even. Hence $\chi$ is a natural DM candidate. Notice that this $\mathbb{Z}_{2}$ symmetry is exact at all loop orders in perturbation theory.
Here we seek a paradigm shift to coherently understand Eq. (4) using the non-invertible Ising fusion rules  [^6] As suggested by the name, conventionally these fusion rules appear in the two-dimensional Ising model, which is one of the prototype non-invertible symmetries from a modern perspective; see e.g. Shao (2023); Schafer-Nameki (2024) for reviews and the references therein. Notice that the fusion rules are identical to that of Tambara-Yamagami (TY) of $\mathbb{Z}_{2}$, i.e. TY$(\mathbb{Z}_{2})$, which in turn could be embedded in e.g. TY$(\mathbb{Z}_{4})$. Even though Eq. (4) respects the fusion rules of TY$(\mathbb{Z}_{4})$, we do not consider such an “unfaithful” scenario.. The algebra is defined for a set consisting of three basis elements $\{\mathbbm{1},\epsilon,\sigma\}$, where $\mathbbm{1}$ denotes the identity element, $\epsilon$ is the element implementing the defining Ising $\mathbb{Z}_{2}$ symmetry, and $\sigma$ is a non-invertible element, meaning that there is no element $x$ such that $\sigma\otimes x=x\otimes\sigma=\mathbbm{1}$. The Ising fusion rules are given by
$\epsilon\otimes\epsilon=\mathbbm{1},\quad\epsilon\otimes\sigma=\sigma\otimes\epsilon=\sigma,\quad\sigma\otimes\sigma=\mathbbm{1}\oplus\epsilon\ ,$ (5)
where $\mathbbm{1}\otimes x=x\otimes\mathbbm{1}=x$ is implicitly understood for $x=\{\mathbbm{1},\epsilon,\sigma\}$. In the DM model of Eq. (4), we label the real scalar $s$ by the element $\epsilon$ and the Majorana fermion $\chi$ by the element $\sigma$, we assume that all the SM fields are labeled by the identity element $\mathbbm{1}$.
Here are the reasons why the Ising fusion rules are useful in understanding the DM model.
  1. 1.
    It is straightforward to check that all the terms in Eq. (4) are consistent with the Ising fusion rules in Eq. (5), hence they are tree-level exact. Furthermore, all the terms (up to dimension four) that are consistent with Eq. (5) are already present in the classical Lagrangian of Eq. (4).
    This explains the peculiar feature that only $s\chi^{2}$ is present in the classical Lagrangian, while $s^{3}$ or $s|H|^{2}$ are not. Naively, $s\chi^{2}$ were forbidden by the Ising $\mathbb{Z}_{2}$ symmetry, but this term is allowed because of the fusion of the non-invertible element in Eq. (5).
  2. 2.
    The all-loop exact $\mathbb{Z}_{2}$ symmetry that guarantees the DM stability follows from the “groupification” of the non-invertible Ising fusion rules Kaidi et al. (2024), which states the following: at arbitrary high loop orders, the non-invertible selection rules reduce to that of a finite Abelian group.  [^7] We refer the readers to Kaidi et al. (2024) for the general discussion of “groupification”. However, in this “calculable” model, we encourage the readers to confirm explicitly the all-loop exactness of the $\mathbb{Z}_{2}$ symmetry protecting the DM stability.
    Notice that the all-loop exact $\mathbb{Z}_{2}$ symmetry is not the Ising $\mathbb{Z}_{2}$ associated with the element $\epsilon$ in Eq. (5). Indeed, the Ising $\mathbb{Z}_{2}$ is violated at the one-loop level in radiative corrections. This is signaled by the one-loop scattering amplitude $s\to ss$, which is induced by closing a loop of $\chi$. (Similarly, the amplitude $s\to H^{\dagger}H$ appears at the two-loop order.)
There are general predictions from the non-invertible Ising fusion rules for the scattering processes for $s$ and $\chi$. For instance, the scatterings of odd numbers of $s$ with itself and any SM particles must be loop-induced (without $\chi$). Indeed, the operators of the type $s^{2n+1}\cdot\mathcal{O}_{\text{SM}}$ for any $n\in\mathbb{Z}_{\geq 0}$ are forbidden by the Ising fusion rules in Eq. (5). On the other hand, the scattering of any number of $s$ with even numbers of $\chi$ can be present at either tree or loop level. This is because the operators of the type $s^{n}\chi^{2m}$ are allowed by the Ising fusion rules for any $n\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$ [^8] If the non-invertible Ising fusion rules were overlooked, one might incorrectly conclude that $s\to\chi\chi$ has to vanish if a $\mathbb{Z}_{2}$ symmetry is imposed under which $s\to-s$. However, as we saw, this process in fact can be at tree level even if the same $\mathbb{Z}_{2}$ is imposed. Again, this shows the power of the non-invertible Ising fusion rules. Of course, the scatterings of $\chi$ are not vanishing only when there are even numbers of them, this is due to the all-loop exact $\mathbb{Z}_{2}$ symmetry. As examples, we sketch the scattering processes $s\to H^{\dagger}H$, $s\to\chi\chi$, and $\chi\chi\to\chi\chi$ in Figure 2. Since all the SM particles are assumed to be labeled by the identity, the Ising fusion rules do not constrain the interactions in $\mathcal{O}_{\text{SM}}$.
Figure 2: A cartoon showing the scattering processes in the DM model with the constraints imposed by the Ising fusion rules. The scattering $s\to H^{\dagger}H$ must be loop induced (denoted by the gray blob), while $s\to\chi\chi$ and $\chi\chi\to\chi\chi$ can be present at tree level (denoted by the white blob) in perturbation theory. These scattering processes are relevant for DM phenomenology and can potentially be tested in experiments.
From the perspective of spurion analysis, the coupling $y_{s\chi}$ needs to be promoted to a background field, and it can be labeled by either $\mathbbm{1}$ or $\epsilon$. On the other hand, the loop-induced coupling $\lambda_{3s}$ of the interaction $s^{3}$ must be labeled by $\epsilon$. From the relation $\lambda_{3s}\sim y_{s\chi}^{3}/(16\pi^{2})$, we conclude that $y_{s\chi}$ should be labeled by $\epsilon$.
We remark that many previous discussions can be carried over to the model where the SM is extended by two gauge-singlet real scalars. The two scalars are denoted as $\epsilon$ and $\sigma$, which are labeled by the corresponding elements in Eq. (5). We consider the following classical Lagrangian,
$\mathcal{L}$ $=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}\epsilon)^{2}+\frac{1}{2}(\partial_{\mu}\sigma)^{2}-\frac{m_{\epsilon}^{2}}{2}\epsilon^{2}-\frac{m_{\sigma}^{2}}{2}\sigma^{2}$
$-\frac{A_{\epsilon\sigma}}{2}\epsilon\,\sigma^{2}-\frac{\lambda_{4\epsilon}}{4!}\epsilon^{4}-\frac{\lambda_{4\sigma}}{4!}\sigma^{4}-\frac{\lambda_{\sigma\epsilon}}{4}\epsilon^{2}\sigma^{2}$
$-\frac{1}{2}(\lambda_{\epsilon h}\epsilon^{2}+\lambda_{\sigma h}\sigma^{2})|H|^{2}\;.~{}$ (6)
The absence of the terms $\epsilon|H|^{2},\sigma|H|^{2},\sigma^{3},\epsilon^{3},\epsilon^{3}\sigma,\sigma^{3}\epsilon$ in contrast to the presence of other terms in classical Lagrangian are elegantly explained using the non-invertible Ising fusion rules. Furthermore, there is the all-loop exact $\mathbb{Z}_{2}$ following from the “groupification” under which $\sigma\to-\sigma$ while all the other particles are charged even. (Again, this all-loop exact $\mathbb{Z}_{2}$ is not to be confused with the Ising $\mathbb{Z}_{2}$ associated with $\epsilon$, where the latter is broken at one loop in perturbation theory.) As a result, $\sigma$ is the DM candidate.
IV Discussion and outlook
In this paper, we demonstrate the power of a general class of non-invertible selection rules in simple particle physics models, where both the Fibonacci and Ising fusion rules can be viewed as some modifications of a $\mathbbm{Z}_{2}$ symmetry. Despite their simplicity, they have distinct phenomenological implications, particularly in particle scatterings.
We see several directions to extend our results.
  • We would like to understand in full generality the spurion analysis for this class of non-invertible fusion rules, perhaps following the inductive proof in Kaidi et al. (2024). A crucial difference is that the couplings, once promoted to background fields, are not propagating degrees of freedom, hence they cannot be glued together.
  • It would be interesting to systematically explore the parameter space of the “non-invertible Ising DM” scenario as a benchmark, following the analysis in Krnjaic (2016); Matsumoto et al. (2019); Chen et al. (2024); Krnjaic (2018). Since the Ising fusion rules impose a distinct pattern of various couplings, the resulting DM phenomenology may exhibit qualitatively novel features. (Similarly, collider studies of the Fibonacci fusion rules in the real scalar model present another intriguing direction.) More generally, the unique scalar interactions predicted by non-invertible selection rules may have profound implications for cosmological phase transitions and gravitational wave experiments.
  • Broadly, we expect nontrivial interplays between the non-invertible selection rules with the more conventional ones, probably from the helicities of the spinning particles, gauge invariance, and ordinary symmetries.
Our results may hint at a paradigm shift, and we hope that they will inspire more explorations into the realm of non-invertible particle physics.
V Acknowledgment
We would like to thank Hao Y. Zhang for bringing Kaidi et al. (2024) to our attention and for the interesting conversation, Di Zhang for helpful comments on the texture zeros and flavor models, Rajath Radhakrishnan for helpful clarifications on fusion rings and categories. M.S. is supported by the MUR projects 2017L5W2PT. M.S. also acknowledges the European Union - NextGenerationEU, in the framework of the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W). The work of L.X.X. is partially supported by European Research Council (ERC) grant n.101039756. L.X.X. would like to thank ITP-CAS for their warm hospitality at the early stage of this work.