### v2

# Phenomenological implications of a class of non-invertible selection rules

Abstract
Through well-motivated models in particle physics, we demonstrate the power of a general class of selection rules arising from non-invertible fusion algebras that are only exact at low orders in perturbation theory. Surprisingly, these non-invertible selection rules can even be applied to the minimal extension of the Standard Model, which is to add a gauge-singlet real scalar. In this model, we show that Fibonacci fusion rules lead to experimentally testable features for the scattering processes of the real scalar. We anticipate that this class of non-invertible selection rules can be applied to a wide range of models beyond the Standard Model. To further strengthen our methodology, we discuss a dark matter model based on the Ising fusion rules, where the dark matter is labeled by the non-invertible element in the algebra, hence its stability is preserved at all loop orders.
I Introduction
The notion of global symmetries has been generalized significantly in the past decade since the seminal work Gaiotto et al. (2015). Perhaps the most drastic generalization is the so-called non-invertible symmetries; see e.g. Cordova et al. (2022); Shao (2023); Schafer-Nameki (2024); Costa et al. (2024); Brennan and Hong (2023) for reviews. Conventionally, these symmetries appear in two-dimensional rational conformal field theories Chang et al. (2019); Thorngren and Wang (2024a, b), and their generalizations in four-dimensional quantum field theories Choi et al. (2022a); Kaidi et al. (2022) and applications in particle physics Choi et al. (2022b); Cordova and Ohmori (2023); Putrov and Wang (2024); Choi et al. (2024); Cordova et al. (2024a, b, c); Delgado and Koren (2025); Kobayashi and Otsuka (2024); Kobayashi et al. (2024); Funakoshi et al. (2024); Kobayashi et al. (2025); Cao et al. (2024) have just been uncovered in recent years. For the interest of particle physics, celebrated examples of non-invertible symmetries include the Adler-Bell-Jackiw anomaly in quantum electrodynamics Choi et al. (2022b); Cordova and Ohmori (2023) and the higher-form generalizations in axion-Maxwell theory Choi et al. (2023); Yokokura (2022); Hidaka et al. (2024); Del Zotto et al. (2024).
More broadly, the existence of non-invertible symmetries suggests an intriguing possibility that the selection rules, when they act on particles, do not have to obey group laws. In this paper, we are interested in a general class of selection rules induced by non-invertible fusion algebras that are only exact in low orders in perturbation theory Kaidi et al. (2024) (see also Heckman et al. (2024)). We call them non-invertible selection rules (NISRs). We aim to show their usefulness in uncovering some peculiar features in particle physics models, which would otherwise be difficult to explain using the selection rules imposed by ordinary symmetries. Surprisingly, these NISRs are even useful in the simplest extensions of the Standard Model (SM). Throughout the paper, we assume that the new particles are weakly coupled, hence perturbation theory can apply.
Let us consider a finite set $A=\{\mathbbm{1},x,y,z,\cdots\}$ of basis elements with the fusion rules $x\otimes y=\sum_{z\in A}N^{z}_{x,y}z$, where $\mathbbm{1}$ is the identity element and $N^{z}_{x,y}$ are the structure constants. In general, the fusion rules are non-invertible. We will only consider the commutative fusion rules, i.e. $x\otimes y=y\otimes x$. One can construct a theory as follows Kaidi et al. (2024):
  • Each field in the theory $\phi_{i}$ is labeled by an element in the set, whose multiplications obey the fusion rules.
  • We assume the existence of the other element which labels the conjugated field $\bar{\phi}_{i}$, such that the identity element $\mathbbm{1}$ exists in the fusion product $\phi_{i}\otimes\bar{\phi}_{i}$.
  • A tree-level interaction term in the classical Lagrangian is allowed when the identity element is contained in the fusion product of the corresponding particles in the interaction.
As such, one constructs all the possible operators consistent with the fusion rules, and they correspond to the tree-level scattering amplitudes. Notice that the kinetic terms of $\phi_{i}$ are always allowed in the above construction. This needs to be contrasted with the background fields which do not propagate, e.g., one can promote coupling constants into the background fields in spurion analysis.
On the other hand, the operators that violate this class of NISRs can be generated in radiative corrections as the loop order is increased. Hence, the NISRs are only exact at low orders in perturbation theory Kaidi et al. (2024), at least tree-level exact. Naively, this contradicts the intuition of spurion analysis for the selection rules imposed by ordinary symmetries, where the selection rules cannot be violated at the loop level in perturbation theory if they are exact at the tree level. However, due to the non-invertible nature, this kind of intuition is invalid for the selection rules discussed in this paper.
In the rest of the paper, we will demonstrate these features in two models in particle physics, arguably the simplest extensions of the SM.
II A real scalar and Fibonacci fusion rules
Let us start by adding a gauge-singlet real scalar to the SM. Despite its simplicity, the model is well-motivated by the rich phenomenology O’Connell et al. (2007); Chen et al. (2015); Buttazzo et al. (2015); He and Zhu (2017); Cao et al. (2018); Corbett et al. (2018); Li et al. (2019); Haisch et al. (2020); Adhikari et al. (2021); Dawson et al. (2021); Lewis et al. (2024). We denote the new scalar particle as $s$, it can have self-interactions and interactions with the Higgs doublet $H$. In particular, we consider the following Lagrangian at the classical level,
$\mathcal{L}=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}s)^{2}-\frac{m_{s}^{2}}{2}s^{2}-\frac{\lambda_{3}}{3!}s^{3}-\frac{\lambda_{4}}{4!}s^{4}-\frac{\lambda_{sh}}{2}s^{2}|H|^{2}\;,$ (1)
where we assume $m_{s}^{2},\lambda_{4},\lambda_{sh}>0$ for simplicity.
Such a Lagrangian is an unconventional one from the perspective of the ordinary $\mathbb{Z}_{2}$ symmetry, under which $s\to-s$ while all the SM particles are uncharged. If the $\mathbb{Z}_{2}$ symmetry is imposed, all the terms with odd powers of $s$ are forbidden, and the $\mathbb{Z}_{2}$ symmetry is exact at all loop orders in perturbation theory. However, the existence of $\lambda_{3}s^{3}$ implies that the $\mathbb{Z}_{2}$ symmetry is explicitly broken, hence there should also be terms such as $\lambda_{1}s$, $\lambda^{\prime}_{sh}s|H|^{2}$ in Eq. (1), if the theory is not fine-tuned. Since there is no obvious “symmetry” that can explain the coupling structure in Eq. (1), it is considered unnatural  [^1] Roughly speaking, the usual naturalness criterion states that a parameter in the Lagrangian is naturally small if there is an enhanced symmetry in the limit when the parameter vanishes ’t Hooft (1980). that $\lambda_{3}s^{3}$ is present in the classical Lagrangian while $\lambda_{1}s$ or $\lambda^{\prime}_{sh}s|H|^{2}$ are not. However, the couplings $\lambda_{1},\lambda^{\prime}_{sh}$ are generated in radiative corrections. For instance, the three-point vertex corresponding to the coupling $\lambda^{\prime}_{sh}$ is generated at one-loop level when $\lambda_{3}$ and $\lambda_{sh}$ are present at tree level:
$(\lambda^{\prime}_{sh})^{\text{1-loop}}\sim\frac{\lambda^{\text{tree}}_{3}\lambda^{\text{tree}}_{sh}}{16\pi^{2}}\ .$ (2)
This equation can be interpreted as the relation between the scattering amplitudes of the processes $\to H^{\dagger}H, ss$$\to ss$, and $ss\to H^{\dagger}H$  [^2] Notice that some particles have to be off-shell due to kinematics. However, our focus here is on the dynamics, namely the interactions between particles. The same consideration applies to the models in the next section., where the first amplitude is obtained by gluing together the legs of $s$ in the second and the third amplitudes.
Based on the above discussion, a natural question arises:
Is there a “symmetry” (or a set of selection rules) that allows the non-vanishing couplings $\lambda_{3},\lambda_{sh},m_{s}^{2},\lambda_{4}$ but forbids $\lambda_{1},\lambda^{\prime}_{sh}$ in the classical Lagrangian?
If such a pattern is realized by a “symmetry”, we understand that the couplings $\lambda_{1},\lambda^{\prime}_{sh}$ are naturally suppressed by a loop factor compared to other tree couplings.
We propose to use the class of NISRs introduced in Kaidi et al. (2024) to address the above question. One peculiar feature is that these selection rules are tree-level exact but will be violated at loop orders in radiative corrections. Hence, they are useful in explaining the hierarchies between the couplings whose relative suppression is given by loop factors. Specifically, we use the Fibonacci fusion rules  [^3] Conventionally, the Fibonacci fusion rules are used to describe anyons in condensed matter physics Trebst et al. (2008). Recently, the same algebra also appeared in flavor models explaining the Yukawa texture zeros based on $\mathbb{Z}_{2}$ gauging of a $\mathbb{Z}_{3}$ symmetry Kobayashi et al. (2024); Funakoshi et al. (2024); Kobayashi et al. (2025)., whose algebra is given by
$\mathbbm{1}\otimes\mathbbm{1}=\mathbbm{1},\quad\quad\mathbbm{1}\otimes\tau=\tau\otimes\mathbbm{1}=\tau,\quad\quad\tau\otimes\tau=\mathbbm{1}\oplus\tau\;.$ (3)
Notice that $\tau$ does not have the inverse, hence this algebra is non-invertible.  [^4] Intuitively, this algebra is very similar to an ordinary $\mathbb{Z}_{2}$ symmetry, that is why it can be relevant to this real scalar model. However, the crucial difference is only due to the non-invertible nature of $\tau$. As we will justify, this is crucial to forbid all the tree-level scattering amplitudes linear in the scalar $s$. In the model, we label the real scalar $s$ by the element $\tau$ and all the other SM particles by the identity $\mathbbm{1}$. It is easy to check that all the terms in Eq. (1) are consistent with Fibonacci fusion rules, i.e., the identity element is produced in the fusion product of the corresponding particles in each of the interactions. More importantly, all of the terms (up to dimension four) consistent with Eq. (3) are already included in the classical Lagrangian. Therefore, the Fibonacci fusion rules are tree-level exact, and they forbid the tree scattering process $s\to H^{\dagger}H$ in a fashion consistent with the usual naturalness criterion (rather than fine-tuning). In contrast, the scattering processes $s\to ss$ and $ss\to H^{\dagger}H$ are not forbidden at the tree level.
More generally, the selection rules in Eq. (3) imply that all the single-$s$ scattering processes must be loop induced, while the multi-$s$ scatterings can be present at tree-level; see in Fig. 1. Indeed, the operators of the type $s\cdot\mathcal{O}_{\text{SM}}$ are forbidden by Eq. (3), while the operators of the types $s^{n}\cdot\mathcal{O}_{\text{SM}}$ are consistent with the selection rules in Eq. (3) for any $n>1$ and $n\in\mathbb{Z}$ [^5] Notice that the tadpole term $\lambda_{1}s$ can be radiatively generated, which in turn induces a nonzero vacuum expectation value for $s$, denoted by $\langle s\rangle$. This is analogous to the Coleman-Weinberg potential Coleman and Weinberg (1973) for radiative symmetry breaking. Consequently, operators of the form $s\langle s\rangle^{n-1}\cdot\mathcal{O}_{\text{SM}}$ are induced radiatively from the tree-level operators $s^{n}\cdot\mathcal{O}_{\text{SM}}$. Similar consideration applies to the model in Eq. (4). Here $\mathcal{O}_{\text{SM}}$ denotes any operator consisting of only the SM particles. Since all the SM particles are labeled by the identity element, the Fibonacci fusion rules do not constrain the interactions in $\mathcal{O}_{\text{SM}}$.
Figure 1: A cartoon showing the scattering processes of the real scalar $s$ and the SM particles whose interactions are constrained by the Fibonacci fusion rules as in Eq. (3). All the single-$s$ scatterings must be loop induced (denoted by the gray blob), while the multi-$s$ scatterings are present at either tree or loop levels (denoted by the white blob) in perturbation theory. If the real scalar is discovered by future experiments, one can potentially test the Fibonacci fusion rules in these scattering processes.
So far, we have applied the Fibonacci fusion rules in Eq. (3) to the real scalar model. It follows that the scatterings of the scalar $s$ need to satisfy the relations such as in Eq. (2). A more general interpretation of this result is that all the single-$s$ scattering processes must be loop-induced. It would be ideal to understand Eq. (2) from the perspective of spurion analysis, which requires promoting the coupling constants into background fields labeled by either $\mathbbm{1}$ or $\tau$, such that the Fibonacci fusion rules are formally preserved. In this example, we see that the tree couplings $\lambda_{3},\lambda_{sh}$ can be labeled by either $\mathbbm{1}$ or $\tau$, but the loop-generated coupling $\lambda^{\prime}_{sh}$ must be labeled by $\tau$. Furthermore, to make Eq. (2) consistent with Eq. (3), at least one of the tree couplings $\lambda_{3},\lambda_{sh}$ should be labeled by $\tau$. We note that this feature differs from the case of ordinary symmetries, where all tree couplings are labeled by the identity whenever the symmetry is exact at tree level. A full understanding of the spurion analysis for the NISRs in Eq. (3) is pursued elsewhere Suzuki et al. (2025).
It is somewhat unexpected that the NISRs in Eq. (3) provide a unified description of both the Fibonacci sequence $\{0,1,1,2,3,5,8,13,21,\cdots\}$ — where the numbers count how many times the identity element appears in successive fusion of $\tau$ with itself — and the scattering properties of a real scalar with SM particles, as shown in Fig. 1. In the correspondence, the initial zero in the Fibonacci sequence reflects the absence of single-$s$ scatterings at tree level. We view this as the simplest example showing the power of the NISRs in particle physics.  [^6] It may seem overkill to apply the NISRs to such a simple model. However, we find this example useful to streamline the overall logic. In a similar spirit, analyzing the ’t Hooft anomaly in the solvable quantum mechanics model Gaiotto et al. (2017); Cheng and Seiberg (2023) is very useful.
III Dark matter and Ising fusion rules
Next, we apply the NISRs to a dark matter (DM) model, demonstrating that our methodology can have a wide range of potential applications.
Let us introduce a gauge-singlet real scalar $s$ and a gauge-singlet Majorana fermion $\chi$, whose interactions are described by the following classical Lagrangian:
$\mathcal{L}$ $=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}s)^{2}-\frac{m_{s}^{2}}{2}s^{2}-\frac{\lambda_{s}}{4!}s^{4}-\frac{1}{2}\lambda_{sh}s^{2}|H|^{2}$ (4)
$+i\chi^{\dagger}\bar{\sigma}^{\mu}\partial_{\mu}\chi-\left(\frac{m_{\chi}}{2}\chi^{2}+\frac{y_{s\chi}}{2}s\,\chi^{2}+\text{h.c.}\right)\;,$~
where $m_{s}^{2},\lambda_{s},\lambda_{sh}>0$. In particular, we assume that the terms $s$, $s^{3}$, and $s|H|^{2}$ are not allowed in the classical Lagrangian. It follows that the scattering processes $s\to ss$ and $s\to H^{\dagger}H$ must be loop-induced. For instance, $s\to ss$ can be induced at the one-loop level by a triangle loop of $\chi$ due to the tree amplitude $s\to\chi\chi$. In turn, $s\to H^{\dagger}H$ can be induced at the two-loop level by gluing together the legs of $s$ in the one-loop amplitude $s\to ss$ and the tree amplitude $ss\to H^{\dagger}H$.
Let us analyze the “symmetry” structure of the classical Lagrangian in Eq. (4). In the paradigm of ordinary symmetries, it is difficult to explain the following two features simultaneously.
  1. 1.
    The interactions of $s$ with the SM particles respect a $\mathbb{Z}_{2}$ symmetry where $s\to-s$, but those with $\chi$ do not. For instance, due to the presence of the $s\chi^{2}$ term in Eq. (4), it is not possible to assign a $\mathbb{Z}_{2}$-odd charge to the scalar $s$. Since there is no $\mathbb{Z}_{2}$ symmetry protecting the interactions of $s$, the absence of the terms $s,s^{3},s|H|^{2}$ in Eq. (4) is not justified.
  2. 2.
    However, Eq. (4) enjoys an invertible $\mathbb{Z}_{2}$ symmetry (i.e., fermion parity) under which $\chi\to-\chi$ while all the other particles are charged even. Hence $\chi$ is a natural DM candidate. This $\mathbb{Z}_{2}$ symmetry is exact at all loop orders in perturbation theory.
[…]
$\epsilon\otimes\epsilon=\mathbbm{1},\quad\epsilon\otimes\sigma=\sigma\otimes\epsilon=\sigma,\quad\sigma\otimes\sigma=\mathbbm{1}\oplus\epsilon\ ,$ (5)
where $\mathbbm{1}\otimes x=x\otimes\mathbbm{1}=x$ is implicitly understood for $x=\{\mathbbm{1},\epsilon,\sigma\}$. In the DM model of Eq. (4), we label the real scalar $s$ by the element $\epsilon$ and the Majorana fermion $\chi$ by the element $\sigma$, and all the SM fields are labeled by the identity element $\mathbbm{1}$.
Here are the reasons why the Ising fusion rules are useful in understanding Eq. (4).
  1. 1.
    It is straightforward to check that all the terms in Eq. (4) are consistent with the Ising fusion rules in Eq. (5), hence they are tree-level exact. Furthermore, all the terms (up to dimension four) that are consistent with Eq. (5) are already present in the classical Lagrangian of Eq. (4).
    This explains the peculiar feature that only $s\chi^{2}$ is present in the classical Lagrangian, while $s,s^{3},s|H|^{2}$ are not. Naively, $s\chi^{2}$ would be forbidden by the $\mathbb{Z}_{2}$ symmetry where $s\to-s$. However, this term is allowed once the fusion rule of the non-invertible element in Eq. (5) is taken into account.
  2. 2.
    The all-loop exact $\mathbb{Z}_{2}$ symmetry (fermion parity), which ensures DM stability, arises from “groupification” of the non-invertible Ising fusion rules Kaidi et al. (2024). In particular, groupification states that at sufficiently high loop orders, NISRs reduce to a finite Abelian group that remains exact up to all loop orders.  [^8] For more detailed discussions, see Kaidi et al. (2024). For completeness, we include a brief review in the Appendix.
    This all-loop exact $\mathbb{Z}_{2}$ symmetry should not be confused with the Ising $\mathbb{Z}_{2}$ associated with the element $\epsilon$ in Eq. (5), where $\epsilon\otimes\epsilon=\mathbbm{1}$. The latter is already violated at one loop through radiative corrections, as signaled by the generation of the process $s\to ss$.
There are general implications from Eq. (5) to the scattering processes of $s$ and $\chi$. For instance, without involving $\chi$, the scatterings of odd numbers of $s$ with itself and any SM particles must be loop-induced. Indeed, the operators of the type $s^{2n+1}\cdot\mathcal{O}_{\text{SM}}$ for any $n\in\mathbb{Z}_{\geq 0}$ are forbidden by the Ising fusion rules in Eq. (5). On the other hand, the scattering of any number of $s$ with even numbers of $\chi$ can be present at either tree or loop levels, since the operators of the type $s^{n}\chi^{2m}$ are allowed by Eq. (5) for any $n\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$ [^9] If the non-invertible Ising fusion rules were overlooked, one might incorrectly conclude that $s\to\chi\chi$ has to vanish if a $\mathbb{Z}_{2}$ symmetry is imposed under which $s\to-s$. However, as we see, this process is actually present at the tree level even when the same $\mathbb{Z}_{2}$ is imposed. Again, this shows the power of the non-invertible Ising fusion rules. The scatterings of $\chi$ do not vanish only when there are even numbers of them, this is due to the all-loop exact $\mathbb{Z}_{2}$ symmetry. As examples, we sketch the scattering processes $s\to H^{\dagger}H$, $s\to\chi\chi$, and $\chi\chi\to\chi\chi$ in Fig. 2. Since all the SM particles are assumed to be labeled by the identity, the Ising fusion rules do not constrain the interactions in $\mathcal{O}_{\text{SM}}$.
Figure 2: A cartoon showing the scattering processes in the DM model with the constraints imposed by the Ising fusion rules. The scattering $s\to H^{\dagger}H$ must be loop induced (denoted by the gray blob), while $s\to\chi\chi$ and $\chi\chi\to\chi\chi$ can be present at tree level (denoted by the white blob) in perturbation theory. These scattering processes are relevant for DM phenomenology and can potentially be tested in experiments.
From the perspective of spurion analysis, the coupling $y_{s\chi}$ needs to be promoted to a background field, and it can be labeled by either $\mathbbm{1}$ or $\epsilon$. On the other hand, the loop-induced coupling $\lambda_{3s}$ of the interaction $s^{3}$ must be labeled by $\epsilon$. From the relation $\lambda_{3s}\sim y_{s\chi}^{3}/(16\pi^{2})$, we conclude that $y_{s\chi}$ should be labeled by $\epsilon$. For a systematic discussion, see Suzuki et al. (2025).
We remark that many previous discussions can be carried over to the model where the SM is extended by two gauge-singlet real scalars. The two scalars are denoted as $\epsilon$ and $\sigma$, which are labeled by the corresponding elements in Eq. (5). We consider the following classical Lagrangian,
$\mathcal{L}$ $=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}\epsilon)^{2}+\frac{1}{2}(\partial_{\mu}\sigma)^{2}-\frac{m_{\epsilon}^{2}}{2}\epsilon^{2}-\frac{m_{\sigma}^{2}}{2}\sigma^{2}$
$-\frac{A_{\epsilon\sigma}}{2}\epsilon\,\sigma^{2}-\frac{\lambda_{4\epsilon}}{4!}\epsilon^{4}-\frac{\lambda_{4\sigma}}{4!}\sigma^{4}-\frac{\lambda_{\sigma\epsilon}}{4}\epsilon^{2}\sigma^{2}$
$-\frac{1}{2}(\lambda_{\epsilon h}\epsilon^{2}+\lambda_{\sigma h}\sigma^{2})|H|^{2}\;.$~ (6)
In contrast to the terms in Eq. (6), the absence of $\epsilon,\sigma,\epsilon\sigma,\epsilon|H|^{2},\sigma|H|^{2},\epsilon^{2}\sigma,\sigma^{3},\epsilon^{3},\epsilon^{3}\sigma,\sigma^{3}\epsilon,\epsilon\sigma|H|^{2}$ is elegantly explained using the non-invertible Ising fusion rules. Furthermore, there is the all-loop exact $\mathbb{Z}_{2}$ following from the “groupification” under which $\sigma\to-\sigma$ while all the other particles are charged even. (Again, this all-loop exact $\mathbb{Z}_{2}$ should not be confused with the Ising $\mathbb{Z}_{2}$ associated with $\epsilon$, where the latter is broken at one loop in perturbation theory.) As a result, $\sigma$ is the DM candidate.
IV Discussion and outlook
In this paper, we demonstrate the power of a general class of NISRs in simple particle physics models, where both the Fibonacci and Ising fusion rules can be viewed as some modifications of a $\mathbbm{Z}_{2}$ symmetry. Despite their simplicity, they have distinct phenomenological implications, particularly in particle scatterings.
We see several directions to extend our results.
  • We would like to understand in full generality the spurion analysis for NISRs. Compared to dynamical particles, a crucial difference is that the couplings, once promoted to background fields, are not propagating degrees of freedom. For relevant discussions on NISRs from near-group fusions, see Suzuki et al. (2025).
  • It would be interesting to systematically explore the parameter space of the “non-invertible Ising DM” scenario as a benchmark, following the analysis in Krnjaic (2016); Matsumoto et al. (2019); Chen et al. (2024); Krnjaic (2018). The Ising fusion rules impose a distinctive pattern of couplings, potentially leading to qualitatively new DM phenomenology. Likewise, collider studies of the Fibonacci fusion rules in the real scalar model present another intriguing direction. More broadly, the unique scalar interactions dictated by NISRs may have profound implications for cosmological phase transitions and gravitational wave signatures.
  • We expect nontrivial interplay between NISRs and conventional selection rules arising from particle helicities, gauge invariance, and ordinary global symmetries.
Our results may hint at a paradigm shift, and we hope that they will stimulate further exploration into the realm of non-invertible particle physics.
V Acknowledgment
We would like to thank Hao Y. Zhang for bringing Kaidi et al. (2024) to our attention and collaboration on the related work on spurion analysis Suzuki et al. (2025), Di Zhang for helpful comments on the texture zeros and flavor models, Rajath Radhakrishnan for helpful clarifications on fusion rings and categories. M.S. is supported by the MUR projects 2017L5W2PT. M.S. also acknowledges the European Union - NextGenerationEU, in the framework of the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W). The work of L.X.X. is partially supported by European Research Council (ERC) grant n.101039756. L.X.X. would like to thank ITP-CAS for their warm hospitality at the early stage of this work.
SUPPLEMENTARY MATERIAL
“Phenomenological implications of a class of non-invertible selection rules”
In this supplementary material, we provide a concise review on the idea of groupification Kaidi et al. (2024), which is crucial for understanding the radiative breaking of NISRs. Along the way, we offer an intuitive way of visualizing it, which facilitates practical applications to particle phenomenology and further motivates the connection with spurion analysis Suzuki et al. (2025).
Appendix A A lightning review on loop-induced groupification of NISRs
It was recently pointed out that NISRs, when applied in perturbation theory, get increasingly violated as the loop order is increased and eventually reduce to those of a finite Abelian group at sufficiently high loop orders. The entire procedure is called the groupification of the original NISRs.
The radiative violation of NISRs (and hence the process of groupification) occurs because some basis elements become indistinguishable at loop orders. More specifically, let $x$ and $y$ be two different basis elements of the set $A$, each labeling at least one particle in the theory. We say that $x$ and $y$ cannot be distinguished at a given loop order $L$ if there exists some element $\omega$ constructed from the products of the conjugate pairs such that:
$x\prec\omega\otimes y\ ,$ (A1)
$\text{where}\quad\omega\prec\bigotimes_{i=1}^{L}z_{i},~z_{i}\prec u\otimes\bar{u}\quad\text{for some $u\in A$}\;.$ (A2)
Here, $\prec$ means that the fusion product on the right-hand side contains the element on the left-hand side. At sufficiently high loop orders (i.e., taking the limit $L\to\infty$), the set made from the elements $w$ stabilizes, and accordingly, the fusion algebra reduces to a finite Abelian group Kaidi et al. (2024).
Figure A1: Diagrammatic illustration of Eqs. (A1) and (A2), where the two particles labeled by $x$ and $y$ become indistinguishable at the one-loop level due to quantum correction.
Eqs. (A1) and (A2) can also be understood diagrammatically. For instance, as shown in Fig. A1, the particles labeled by $x$ and $y$ are indistinguishable at the one-loop order — namely $L=1$ — if there exists an element $z\prec u\otimes\bar{u}$ such that $y\prec z\otimes x$. (However, notice that $x$ and $y$ are distinct at the tree level, where the two-point vertex mixing $x$ and $y$ is forbidden.) In a sequence, the diagrams in Fig. A1 are as follows:
  1. 1.
    Due to quantum correction, the particle labeled by $x$ can split into two virtual particles labeled by $u$ and $v$, respectively. Here, the elements satisfy the relation $v\prec x\otimes u$ in the fusion algebra.
  2. 2.
    We cut open the internal line for the virtual particle labeled by $u$. The original one-loop diagram becomes a tree-level one, where the elements from the left to right satisfy the relations $v\prec x\otimes u,\ y\prec v\otimes\bar{u}$ in the fusion algebra. From these relations, it is easy to see that $y\prec x\otimes u\otimes\bar{u}$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed, i.e., $z\prec u\otimes\bar{u},\ y\prec x\otimes z$. Notice that $z$ does not have to correspond to a dynamical particle in the theory, but it may be related to the label $v$ through the associators (i.e., the $F$-symbols). A systematic understanding is left for future work.
  4. 4.
    Finally, we glue together the conjugate pair $u$ and $\bar{u}$. We see that the elements $x$ and $y$ become indistinguishable by dressing the element $z$, which in turn is produced by $u\otimes\bar{u}$ at the one-loop order.
The same reasoning applies to higher loop orders, where one can always cut open the internal lines, such that the loop diagram reduces to a tree-level one, whose external legs contain the original external lines plus some number of additional conjugate pairs. Here, the number of additional conjugate pairs matches the internal lines being cut. After suitable deformation, one can glue the conjugate pairs together, and then one can explicitly see that some elements become indistinguishable precisely because of the dressing of conjugate pairs, i.e., Eqs. (A1) and (A2).
From the perspective of the spurion analysis, one can absorb the elements $z_{i}$ produced by the conjugate pairs $u\otimes\bar{u}$, which label dynamical particles, into the definition of coupling constants. (Therefore, one can choose to replace all conjugate pairs with the identity element in the spurion analysis.) For related discussions of NISRs from near-group fusion algebras, see Suzuki et al. (2025).
In the following, we apply the above methodology to the models studied in the main text.
Appendix B Example: A real scalar and Fibonacci fusion rules
For the NISRs from the Fibonacci fusion algebra, the process of groupification is completed already at the one-loop order. Namely, the set of elements $w$ produced by the conjugate pairs stabilizes at the one-loop order, i.e., $\{w\}=\{\mathbbm{1},\tau\}$ at $L=1$ where both elements are obtained from the fusion product $\tau\otimes\tau$; see Eq. (A2). By dressing $\tau\in\{w\}$, we find $\mathbbm{1}\prec\tau\otimes\tau$ as in Eq. (A1), i.e., the two elements $\mathbbm{1}$ and $\tau$ become indistinguishable and equivalent at the one-loop order. As a result, the Fibonacci fusion algebra reduces to the trivial group, which consists only of the identity as its group element.
In the real gauge-singlet scalar model, the argument of groupification can be visualized by the diagrams shown in Fig. B1. Specifically, we consider the three-point vertex $s|H|^{2}$, where the real scalar $s$ is labeled by the element $\tau$ while the Higgs doublet $H$ (hence $H^{\dagger}H$) is labeled by $\mathbbm{1}$. Clearly, such a vertex is forbidden at the tree level, but it is generated at the one-loop order. Correspondingly, the relevant process $s\to|H|^{2}$ can occur precisely because $\tau$ and $\mathbbm{1}$ become indistinguishable and equivalent. In a sequence, the diagrams in Fig. B1 are as follows:
  1. 1.
    Due to quantum correction, the real scalar $s$ can split into two virtual scalars $s$. Here, the corresponding elements satisfy the relation $\tau\prec\tau\otimes\tau$ in the Fibonacci fusion algebra.
  2. 2.
    We cut open the internal line for $s$. The original one-loop diagram becomes a tree-level one, where the elements from left to right satisfy the relations $\tau\prec\tau\otimes\tau,\ \mathbbm{1}\prec\tau\otimes\tau$ in the Fibonacci fusion algebra. From these relations, it is easy to see that $\mathbbm{1}\prec\tau\otimes\tau\otimes\tau$. Notice that $\tau$ is self-conjugate, namely $\bar{\tau}=\tau$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed.
  4. 4.
    Finally, we glue together the conjugate pair back to restore the loop order. We see that the elements $\tau$ and $\mathbbm{1}$ become indistinguishable by dressing the element $\tau$, which in turn is produced by $\tau\otimes\tau$ at the one-loop order.
Figure B1: Diagrammatic illustration of groupification for the Fibonacci fusion algebra, which is concretely realized in the real scalar model. Specifically, here we consider the three-point vertex $s|H|^{2}$.
From the viewpoint of the spurion analysis, we can choose to absorb the element $\tau$ in the last step — which is produced by the fusion of a conjugate pair $\tau\otimes\tau$ — into the definition of the coupling constant for the three-point vertex $s|H|^{2}$, i.e., $\lambda^{\prime}_{sh}\sim\tau$. In turn, it is related to the tree couplings $\lambda_{3}$ and $\lambda_{sh}$ respectively for the vertices $s^{3}$ and $s^{2}|H|^{2}$. Following Suzuki et al. (2025), we label the tree couplings as $\lambda_{3}\sim\tau$ and $\lambda_{sh}\sim\mathbbm{1}$, such that the relation between the couplings $\lambda^{\prime}_{sh}\sim\lambda_{3}\lambda_{sh}/(16\pi^{2})$ is consistent with the fusion rule $\tau\prec\tau\otimes\tau$.
Appendix C Example: Dark matter and Ising fusion rules
For the NISRs from the Ising fusion algebra, the groupification is again completed at the one-loop order, since the set of elements $w$ produced by the conjugate pairs stabilizes at the one-loop order, i.e., $\{w\}=\{\mathbbm{1},\epsilon\}$ at $L=1$, where $\mathbbm{1}\prec\epsilon\otimes\epsilon$, $\mathbbm{1}\prec\sigma\otimes\sigma$, and $\epsilon\prec\sigma\otimes\sigma$. These equations are in the same form as Eq. (A2). Furthermore, by dressing with $\epsilon$, we find that $\mathbbm{1}\prec\epsilon\otimes\epsilon$ and $\sigma\prec\sigma\otimes\epsilon$, which implies that at the one-loop order $\epsilon$ becomes indistinguishable and equivalent to $\mathbbm{1}$, while $\sigma$ remains distinct; see Eq. (A1). Consequently, at the one-loop order and higher, the Ising fusion algebra reduces to a $\mathbb{Z}_{2}$ group, where the $\mathbb{Z}_{2}$-odd charge is generated by $\sigma$.
In the dark matter model, the argument of groupification can be visualized by the diagrams shown in Fig. C1. Specifically, we consider the three-point vertex $s^{3}$ that is forbidden at the tree level but is generated at the one-loop order. In the Ising fusion algebra, the mediator $s$ and the dark matter $\chi$ are labeled by the elements $\epsilon$ and $\sigma$, respectively. Hence, the process $s\to ss$ can occur precisely because $\epsilon$ and $\mathbbm{1}$ become indistinguishable and equivalent at the one-loop order. In a sequence, the diagrams in Fig. C1 are as follows:
  1. 1.
    Due to quantum correction, the $s$ particle can split into two virtual $\chi$ particles. Here, the corresponding elements satisfy the relation $\sigma\prec\sigma\otimes\epsilon$ in the Ising fusion algebra.
  2. 2.
    We cut open the internal line for $\chi$. The original one-loop diagram becomes a tree-level one, where the elements from left to right satisfy the relations $\sigma\prec\epsilon\otimes\sigma,\ \sigma\prec\epsilon\otimes\sigma,\ \epsilon\prec\sigma\otimes\sigma$ in the Ising fusion algebra. From these relations, it is easy to see that $\epsilon\prec\epsilon\otimes\sigma\otimes\epsilon\otimes\sigma$. Notice that $\sigma$ is self-conjugate, namely $\bar{\sigma}=\sigma$.
  3. 3.
    We deform the tree diagram obtained in the last step, such that the order of fusions is changed. The corresponding elements from left to right satisfy the relations $\epsilon\prec\epsilon\otimes\epsilon\otimes\epsilon$, $\sigma\prec\sigma\otimes\epsilon$.
  4. 4.
    Finally, we glue together the conjugate pair back to restore the loop order. We see that the elements $\epsilon$ and $\mathbbm{1}=\epsilon\otimes\epsilon$ become indistinguishable by dressing the element $\epsilon$, which in turn is produced by $\sigma\otimes\sigma$ at the one-loop order.
Figure C1: Diagrammatic illustration of groupification for the Ising fusion algebra, which is concretely realized in the dark matter model. Specifically, we consider the three-point vertex $s^{3}$ here.
From the viewpoint of the spurion analysis, we can choose to absorb the element $\epsilon$ in the last step — which is produced by the fusion of the conjugate pair $\sigma\otimes\sigma$ — into the definition of the coupling constant for the three-point vertex $s^{3}$, denoted by $\lambda_{3s}\sim\epsilon$. In turn, it is related to the tree coupling $y_{s\chi}$ for the vertex $s\chi^{2}$. Following Suzuki et al. (2025), we label the tree couplings as $y_{s\chi}\sim\epsilon$, such that the relation between the couplings $\lambda_{3s}\sim y^{3}_{s\chi}/(16\pi^{2})$ is consistent with the fusion rule $\epsilon\prec\epsilon\otimes\epsilon\otimes\epsilon$.